White light containing wavelengths from 4.10E2 nm to7.50E2 nm falls on a grating with 9.2E3 lines/cm. How wide is thefirst-order spectrum on a screen 2.50 m away?
Answer #1
first get d
       d = 1/N =   1 / 9200 lines per cm   =
           =     1.086957 x 10-4 cm =    1086.957 nm
Now use
    λ = dsinθ    to get θ for the longest andshortest wavelength
     sinθ = λ /d
for shortest wavelength of 410 nm
     sinθ = 410 /1086.957      θ =   22.16degrees
for longest wavelength of   750 nm
   sinθ = 750 /1086.957        θ =   43.63 degrees
Now get the distance from center for each by usingtangent...
   x = D tanθ
shortest?
   x = 2.50 * tan22.16 =     1.0182 meters
longest?
   x = 2.50 * tan43.63  =   2.3832 meters
width = Δx = 2.3832 - 1.0182 =   1.365 meters